# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v1

@@ Abstract
Although the Harrow-Hassidim-Lloyd (HHL) algorithm offers an exponential speedup in system size for treating linear equations of the form $A\vec{x}=\vec{b}$ on quantum computers when compared to their traditional counterparts, it faces a challenge related to the condition number ($\mathcal{\kappa}$) scaling of the $A$ matrix. In this work, we address the issue by introducing the post-selection-improved HHL (Psi-HHL) approach that involves a simple yet effective modification of the HHL algorithm to extract a feature of |$\ket{x}$, and which leads to achieving optimal behaviour in $\mathcal{\kappa}$ (linear scaling) for large condition number situations. This has the important practical implication of having to use substantially fewer shots relative to the traditional HHL algorithm. We carry out two sets of simulations, where we go up to 26-qubit calculations, to demonstrate the ability of Psi-HHL to handle situations involving large $\mathcal{\kappa}$ matrices via: (a) a set of toy matrices, for which we go up to size $64\times 64$ and $\mathcal{\kappa}$ values of up to $\approx$ 1 million, and (b) a deep-dive into quantum chemistry, where we consider matrices up to size $256\times 256$ that reach $\mathcal{\kappa}$ of about 466. The molecular systems that we consider are Li2, RbH, and CsH. Although the feature of |$\ket{x}$ considered in our examples is an overlap between the input and output states of the HHL algorithm, our approach is general and can be applied in principle to any transition matrix element involving |$\ket{x}$.
@@ I Introduction
Despite possessing the aforementioned advantages, the HHL algorithm relies upon the probability of successfully post-selecting $1$ on the HHL ancilla qubit, that is, $P(1)$ [1]. It is known that there is a substantial decrease in $P(1)$ with increase in the condition number of the problem matrix $A$, which we shall denote by $\mathcal{\kappa}$. In other words, the fraction of usable shots in the HHL algorithm becomes very low when large $\mathcal{\kappa}$ input matrices are involved. Unfortunately, it is not uncommon to encounter the issue while employing HHL for some interesting problems (for example, see Ref. [3] for application of the algorithm to hydrological systems), including quantum chemistry, which we shall discuss in detail in our Results section. Several works in literature have addressed the issue of handling large $\mathcal{\kappa}$ matrices either directly (where $\mathcal{\kappa}$ of the $A$ matrix itself is altered via approaches such as preconditioning) or indirectly [4, 5, 6, 7, 8, 3, 9]. For a thorough survey, we direct the reader to Section 1.2 of Ref. [10].
@@
The paper is organized as follows: Section II presents the theory and methodology involved in our work by introducing the Psi-HHL approach, including a discussion on the complexity theoretic aspects (in $\mathcal{\kappa}$) of Psi-HHL. Section III presents our results for Psi-HHL applied to (a) suitably chosen sets of $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$ toy matrices, as well as for (b) molecular calculations ($64\times 64$ and $256\times 256$ $A$ matrix sizes) in the linearized coupled cluster (LCC) HHL and Psi-HHL frameworks. We conclude in Section IV.
Figure 1: (a) Schematic of the HHL algorithm, where 1 is post-selected and for which $P(1)$ would be low for large $\mathcal{\kappa}$. |$\ket{b}$ is the input state register with $n_{b}$ qubits, while $n_{r}$ refers to the number of clock register qubits. The Hong-Ou-Mandel (HOM) module is a controlled-SWAP test without an additional ancilla qubit, and we use it to extract |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in. Sub-figure (b) provides a block-diagram description of the HHL and the HOM modules. $QPE$ refers to the quantum phase estimation while $CR(\theta)$ refers to the controlled-rotation module that inverts the eigenvalues of $A$. The dashed cambered rectangle around the CNOT gate in the HOM circuit module, and which is accompanied by $(i)$, indicates that the gate is controlled on the ith qubit of the |$\ket{b}$ register. Sub-figure (c) consists of HHL1 in the top panel, where $P(0)$ is post-selected, and HHL2 in the bottom panel, where the traditional HHL module is appended with an additional mixing unitary, $R_{Y}(2\alpha)$, and where $P^{\prime}(1)$ is post-selected. Psi-HHL relies on subtracting the final outcomes of HHL1 and HHL2, and as the expression at the bottom of panel (c) indicates, we recover the correlation energy via the procedure to within a simple factor. In our notation, $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the correlation energies obtained from HHL and Psi-HHL respectively. Further, $a_{i},b_{i},c_{i},d_{i}\in\mathbb{R}$ are constants.
@@ II.1 The HHL algorithm
We briefly sketch the relevant details of the HHL algorithm (we refer the readers to Refs. [14, 15, 16, 17] for detailed descriptions of the algorithm) in Figs. 1(a) and 1(b), where an amplitude-encoded input |$\arrowvert b\rangle=\sum_{i}b_{i}\arrowvert\nu_{i}\rangle$, such that $\sum_{i}|b_{i}|^{2}=1$ where {|$\{\arrowvert\nu_{i}\rangle\}$ refers to the eigenbasis of $A$, is fed into the HHL algorithm along with $n_{r}$ number of clock register qubits and one HHL ancillary qubit. The HHL algorithm itself involves QPE, followed by a non-trivial controlled-rotation module, and then a QPE module. At the end of the circuit, we have the following state vector,
@@
respectively, where subscript ‘$\mathrm{un}$’ refers to unnormalized states— $|b\rangle_{\rm un}$ is the unnormalized input state after amplitude encoding and $|x\rangle_{\rm un}=\sum_{i}\frac{b_{i}C}{\tilde{\lambda}_{i}}|\nu_{i}\rangle$, with $\sum_{i}|b_{i}|^{2}=1$. The probability of successfully post-selecting outcome ‘$1$’ is given as $P(1)=\||x\rangle_{\rm un}\|^{2}$. Finally, when one wants to predict a feature of the solution vector, say, the correlation energy of a molecular system (proportional to the overlap between |$\ket{x}$ and |$\ket{b}$), it can be done by appending an additional circuit (the Hong-Ou-Mandel module [18]), abbreviated as HOM in Fig. 1(b) that inputs the output, |$\arrowvert x\rangle$, of the HHL algorithm and |$\arrowvert b\rangle$, to output the overlap between the two quantities. The correlation energy is then calculated as
@@ II.2 The Psi-HHL approach
In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Fig. 1(c). The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Fig. 1(a) and Fig. 1(b). The HHL2 entails traditional HHL module appended with a mixing $2\times 2$ unitary, given as $R_{Y}(2\alpha)$ and post-selecting 1 from the mixed signal. The role of this unitary is to mix the coefficients of $|0\rangle$ and $|1\rangle$ in the traditional HHL (see Eq. 1), thereby outputting a ‘mixed’ signal. The core idea behind Psi-HHL is to subtract the correlation energies associated with HHL1 and HHL2 thus leaving us with the correlation energy corresponding to the ‘correct’ signal. The gain in taking this route is the efficiency that we achieve in the number of shots expended. We summarize this in the following proposition:
@@ II.2.2 HHL2: Computing the mixed signal
The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Fig. 1(c). This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$. We recall that $R_{Y}(2\alpha)$ is a rotation about the $Y$ axis on the Bloch sphere as given below
@@ II.2.3 Psi-HHL: Subtracting signals
By comparing the right hand sides of Eq. II.1 and Eq. 14, we see that the expressions for $E_{\mathrm{c}}^{\rm HHL}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ differ only by a factor of $\rm cot(\alpha)$ as shown below.
@@ II.3 Complexity aspects
where $\mathbb{I}$ is the identity operator written in the eigenbasis of operator $A$, constant $C=O(1/\mathcal{\kappa})$, where the condition number $\mathcal{\kappa}$ is taken as $\lambda_{\rm max}/\lambda_{\rm min}$ and $|b\rangle$ is the input to the HHL circuit. The set of values $\bigg{\{}\sqrt{1-\frac{C^{2}}{\lambda_{i}^{2}}}\bigg{\}}$ are the eigenvalues of the operator $\sqrt{\mathbb{I}-C^{2}A^{-2}}$, with the largest eigenvalue being $\sqrt{1-\frac{C^{2}}{\lambda_{\rm max}^{2}}}$, and therefore
@@
The first line uses the property of sub-multiplicativity of matrix norms, that is, $\|AB\|\leq\|A\|\|B\|$. In Eq. II.3, we recall that we set $C=\lambda_{\rm min}$, .$\|.\|_{\rm op}$ is the operator norm, and $\||b\rangle\|=1$. Note that the smallest eigenvalue of this operator is $\sqrt{1-\frac{C^{2}}{\lambda_{\rm min}^{2}}}=0$, and corresponds to $P(0)=0$, which leads to expending a large number of shots in HHL1. One could circumvent this issue by appropriately choosing a scaled $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, leading to $P(0)\geq 1-\gamma^{2}$. This idea has been illustrated via an example as shown in Fig. 2, where we choose the scaling $\gamma=0.5$.
@@
$P^{\prime}(1)$$\approx$ $\rm\|[\sqrt{(\mathbb{I}-C^{2}A^{-2}})\mathrm{sin}(\alpha)$ (18)
@@
$P^{\prime}(1)$$\leq$ $\rm\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
@@
For the choice of $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, one can push the lower bound for the choice of given $\alpha$. This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Fig. 1(a). We show this in Fig. 2, where for a given $\alpha=60^{\circ}$, we obtain

### v2

@@ Abstract
Although the Harrow-Hassidim-Lloyd (HHL) algorithm offers an exponential speedup in system size for treating linear equations of the form $A\vec{x}=\vec{b}$ on quantum computers when compared to their traditional counterparts, it faces a challenge related to the condition number ($\mathcal{\kappa}$) scaling of the $A$ matrix. In this work, we address the issue by introducing the post-selection-improved HHL (Psi-HHL) approach that involves a simple yet effective modification of the HHL algorithm to extract a feature of |$\ket{x}$, and which leads to achieving optimal behaviour in $\mathcal{\kappa}$ (linear scaling) for large condition number situations. This has the important practical implication of having to use substantially fewer shots relative to the traditional HHL algorithm. Additionally, we discuss the capability of Psi-HHL to address the cases where $A$ is singular. We carry out two sets of simulations, where we go up to 26-qubit calculations, to demonstrate the ability of Psi-HHL to handle situations involving large $\mathcal{\kappa}$ matrices via: (a) a set of toy matrices, for which we go up to size $64\times 64$ and $\mathcal{\kappa}$ values of up to $\approx$ 1 million, and (b) application to quantum chemistry, where we consider matrices up to size $256\times 256$ that reach $\mathcal{\kappa}$ of about 393. The molecular systems that we consider are Li2, KH, RbH, and CsH. Although the feature of |$\ket{x}$ considered in our examples is an overlap between the input and output states of the HHL algorithm, our approach is general and can be applied in principle to any transition matrix element involving |$\ket{x}$.
@@ I Introduction
Despite possessing the aforementioned advantages, the HHL algorithm relies upon the probability of successfully post-selecting $1$ on the HHL ancilla qubit, that is, $P(1)$ [1]. It is known that there can be a substantial decrease in $P(1)$ with increase in the condition number of the problem matrix $A$, which we shall denote by $\mathcal{\kappa}$. In other words, the fraction of usable shots in the HHL algorithm becomes very low when large $\mathcal{\kappa}$ input matrices are involved. Unfortunately, it is not uncommon to encounter the issue while employing HHL for some interesting problems (for example, see Ref. [3] for application of the algorithm to hydrological systems), including quantum chemistry, which we shall discuss in detail in our Results section. Several works in literature have addressed the issue of handling large $\mathcal{\kappa}$ matrices either directly (where $\mathcal{\kappa}$ of the $A$ matrix itself is altered via approaches such as preconditioning) or indirectly [4, 5, 6, 7, 8, 3, 9]. For a thorough survey, we direct the reader to Section 1.2 of Ref. [10].
@@
The paper is organized as follows: Section II presents the theory and methodology involved in our work by introducing the Psi-HHL approach, including a discussion on the complexity theoretic aspects (in $\mathcal{\kappa}$) of Psi-HHL and on the ability of the algorithm to handle instances where $A$ is singular. Section III presents our results for Psi-HHL applied to (a) suitably chosen sets of $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$ toy matrices, as well as for (b) molecular calculations ($64\times 64$ and $256\times 256$ $A$ matrix sizes) in the linearized coupled cluster (LCC) HHL and Psi-HHL frameworks. We conclude in Section IV.
Figure 1: (a) Schematic of the HHL algorithm, where 1 is post-selected and for which $P(1)$ would be low for large $\mathcal{\kappa}$. |$\ket{b}$ is the input state register with $n_{b}$ qubits, while $n_{r}$ refers to the number of clock register qubits. The Hong-Ou-Mandel (HOM) module is a controlled-SWAP test without an additional ancilla qubit, and we use it to extract |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in. Sub-figure (b) provides a block-diagram description of the HHL and the HOM modules. $QPE$ refers to the quantum phase estimation while $CR(\theta)$ refers to the controlled-rotation module that inverts the eigenvalues of $A$. The $CR(\theta)$ circuit is built out of uniformly controlled rotations involving a set of angles $\{\theta_{i}\}$, denoted compactly as just $\theta$ in the figure. The dashed cambered rectangle around the CNOT gate in the HOM circuit module, and which is accompanied by $(i)$, indicates that the gate is controlled on the ith qubit of the |$\ket{b}$ register. Sub-figure (c) consists of HHL1 in the top panel, where $P(0)$ is post-selected, and HHL2 in the bottom panel, where the traditional HHL module is appended with an additional mixing unitary, $R_{Y}(2\alpha)$, and where $P^{\prime}(1)$ is post-selected. Psi-HHL relies on subtracting the final outcomes of HHL1 and HHL2, and as the expression at the bottom of panel (c) indicates, we recover the correlation energy via the procedure to within a simple factor. In our notation, $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the correlation energies obtained from HHL and Psi-HHL respectively. Further, $a_{i},b_{i},c_{i},d_{i}\in\mathbb{R}$ are constants.
@@ II.1 The HHL algorithm
We briefly sketch the relevant details of the HHL algorithm (we refer the readers to Refs. [14, 15, 16, 17] for detailed descriptions of the algorithm) in Figures 1(a) and 1(b), where an amplitude-encoded input |$\arrowvert b\rangle=\sum_{i}b_{i}\arrowvert\nu_{i}\rangle$, such that $\sum_{i}|b_{i}|^{2}=1$ where {|$\{\arrowvert\nu_{i}\rangle\}$ refers to the eigenbasis of $A$, is fed into the HHL algorithm along with $n_{r}$ number of clock register qubits and one HHL ancillary qubit. The HHL algorithm itself involves QPE, followed by a non-trivial controlled-rotation module, and then a QPE module. At the end of the circuit, we have the following state vector,
@@
respectively, where subscript ‘$\mathrm{un}$’ refers to unnormalized states— $|b\rangle_{\rm un}$ is the unnormalized input state after amplitude encoding and $|x\rangle_{\rm un}=\sum_{i}\frac{b_{i}C}{\tilde{\lambda}_{i}}|\nu_{i}\rangle$, with $\sum_{i}|b_{i}|^{2}=1$. The probability of successfully post-selecting outcome ‘$1$’ is given as $P(1)=\||x\rangle_{\rm un}\|^{2}$. Finally, when one wants to predict a feature of the solution vector, say, the correlation energy of a molecular system (proportional to the overlap between |$\ket{x}$ and |$\ket{b}$), it can be done by appending an additional circuit (the Hong-Ou-Mandel module [18]), abbreviated as HOM in Figure 1(b) that inputs the output, |$\arrowvert x\rangle$, of the HHL algorithm and |$\arrowvert b\rangle$, to output the overlap between the two quantities. The correlation energy is then calculated as
@@ II.2 The Psi-HHL approach
In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Figure 1(c). The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Figure 1(a) and Figure 1(b). The HHL2 entails traditional HHL module appended with a mixing $2\times 2$ unitary, given as $R_{Y}(2\alpha)$ and post-selecting 1 from the mixed signal. The role of this unitary is to mix the coefficients of $|0\rangle$ and $|1\rangle$ in the traditional HHL (see Eq. 1), thereby outputting a ‘mixed’ signal. The core idea behind Psi-HHL is to subtract the correlation energies associated with HHL1 and HHL2 thus leaving us with the correlation energy corresponding to the ‘correct’ signal. The gain in taking this route is the efficiency that we achieve in the number of shots expended. We summarize this in the following proposition:
@@ II.2.2 HHL2: Computing the mixed signal
The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Figure 1(c). This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$. We recall that $R_{Y}(2\alpha)$ is a rotation about the $Y$ axis on the Bloch sphere as given below
@@ II.2.3 Psi-HHL: Subtracting signals
By comparing the right hand sides of Eq. II.1 and Eq. 14, we see that the expressions for $E_{\mathrm{c}}^{\rm HHL}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the same as shown below:
@@ II.3 Complexity aspects
where $\mathbb{I}$ is the identity operator written in the eigenbasis of operator $A$, constant $C=\mathcal{O}(1/\mathcal{\kappa})$, where the condition number $\mathcal{\kappa}$ is taken as $\lambda_{\rm max}/\lambda_{\rm min}$ and $|b\rangle$ is the input to the HHL circuit. The set of values $\bigg{\{}\sqrt{1-\frac{C^{2}}{\lambda_{i}^{2}}}\bigg{\}}$ are the eigenvalues of the operator $\sqrt{\mathbb{I}-C^{2}A^{-2}}$, with the largest eigenvalue being $\sqrt{1-\frac{C^{2}}{\lambda_{\rm max}^{2}}}$, and therefore
@@
The first line uses the property of sub-multiplicativity of matrix norms, that is, $\|AB\|\leq\|A\|\|B\|$. In Eq. II.3, we recall that we set $C=\lambda_{\rm min}$, .$\|.\|_{\rm op}$ is the operator norm, and $\||b\rangle\|=1$. Note that the smallest eigenvalue of this operator is $\sqrt{1-\frac{C^{2}}{\lambda_{\rm min}^{2}}}=0$, and corresponds to $P(0)=0$, which leads to expending a large number of shots in HHL1. One could circumvent this issue by appropriately choosing a scaled $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, leading to $P(0)\geq 1-\gamma^{2}$. This idea has been illustrated via an example as shown in Figure 2, where we choose the scaling $\gamma=0.5$.
@@
$P^{\prime}(1)$$\approx$ $\|[\sqrt{(\mathbb{I}-C^{2}A^{-2}})\mathrm{sin}(\alpha)$ (18)
@@
$P^{\prime}(1)$$\leq$ $\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
@@
For the choice of $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, one can push the lower bound for the choice of given $\alpha$. This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Figure 1(a). We show this in Figure 2, where for a given $\alpha=60^{\circ}$, we obtain
@@
II.4 Singular matrices
When the HHL algorithm is executed for a large system size for some target application, the user ideally should not possess knowledge of some crucial properties of $A$, including its condition number and invertibility. In the extreme event that $A$ happens to be singular and its condition number is not known a priori, the algorithm always outputs 00, that is, $P(1)=0$, and a user has to supply a rather large number of shots before perhaps drawing on the realization that the problem may not produce a $1$. However, Psi-HHL, by construction, circumvents this issue that HHL suffers from. For illustration, we pick two small but extreme examples, as follows,
  • A $2\times 2$ matrix given by $A=\begin{pmatrix}0.25&0.00\\ 0.00&0.00\end{pmatrix}$ and the $\vec{b}$ is $\begin{pmatrix}0.00\\ 1.00\end{pmatrix}$,
  • A $4\times 4$ matrix $A=\begin{pmatrix}1.00&2.00&0.00&0.00\\ 2.00&4.00&0.00&0.00\\ 0.00&0.00&0.00&0.00\\ 0.00&0.00&0.00&2.00\end{pmatrix}$ and with $\vec{b}=\begin{pmatrix}0.00\\ 0.00\\ 1.00\\ 0.00\end{pmatrix}$.
In both the cases above $\vec{b}\in N(A)$, such that $N(A)$ is the null space of $A$ matrix. We execute HHL as well as Psi-HHL for these two cases by setting $C/\lambda_{i}$=00 whenever the eigenvalues of $A$ are zero, that is, $\lambda_{i}=0$. The traditional HHL algorithm would expend all the supplied shots without being able to post-select outcome $1$, since $P(1)=0$. The HHL1 of Psi-HHL successfully post-selects outcome 00 with probability $P(0)=1$. On the other hand, the HHL2 module of Psi-HHL post-selects outcome $1$ with probability $\sin^{2}(\alpha)$, see Eq. 10. The data points presented in Figure 3 demonstrate these probabilities for the two examples considered. Both these HHLs yield the solution $|x\rangle=|b\rangle$. The overlap $|\langle x|b\rangle|$ from the HOM module in HHL1 and HHL2 both yield a value of 1, with all of the additional $\sin({\alpha})$ factors cancelling out (see Figure 1). Subtracting the two overlap values results in 0, as expected. Thus, the Psi-HHL approach proves powerful when $P(0)\gg P(1)$. We now consider the case when $\vec{b}$ is an eigenvector of $A$. Here too, one could still get the same result via Psi-HHL by choosing the evolution time, $t$, as a multiple of $2\pi$, which could be incorrect. Therefore, in order to reliably calculate the outputs from both the cases using Psi-HHL, one chooses a non-trivial $t$ in the QPE module, such that it is not an integral multiple of $2\pi$, in order to avoid the possibility of obtaining $\lambda_{i}=0$ for both the cases. Choosing a non-trivial $t$ will resolve this issue, in the following way:
  • In the case where $\vec{b}$ is an eigenvector of $A$, one obtains an eigenvalue $\lambda_{i}\neq 0$, yielding an associated nontrivial controlled-rotation angle, $\theta_{i}$.
  • However, when $\vec{b}$ is a null space vector, one would still obtain an eigenvalue $\lambda_{i}=0$ leading to a controlled-rotation angle of 00.
          (a)                             (b)
Figure 3: Figure showing the performance of HHL and Psi-HHL for singular matrices of size $2\times 2$ (sub-figure (a)) and $4\times 4$ (sub-figure (b)).